OpenStudy (anonymous):

According to the commutative property, 7x - 10y is equal to: A. 10y - 7x B. -10y + 7x C. 10x - 7y D. -3xy

OpenStudy (nenadmatematika):

b

hero (hero):

Commutative Property states that a + b = b + a. In this case, 7x - 10 = -10 + 7x